the present age of reshma's father is three times her present age after five years the sum of their age will be 70 find their present age

Answer by PR1375(2)

Let the present age of reshma be X,
Then her fathers age 3X
After 5 Years reshmas age become X+5
Fathers age become 3X+5
Their total age after 5 Years is 70 So,
(X+5)+(3X+5)=70
4X+10=50
4X=70-10
4X=60
X=60/4 =15
X=15
Reshma's age is 15
Father's age is 3×15=45